OpenStudy (anonymous):

identify the sequence as arithmetic, geometric, or neither. explain your answer. 1.6, 0.8, 0.4, 0.2

OpenStudy (anonymous):

@paki

hartnn (hartnn):

do u see any pattern in that sequence ?

hartnn (hartnn):

hint : if you find a common ratio, its geometric sequence if you find any common difference, its arithmetic sequence else neither

OpenStudy (anonymous):

um, arithmetric

hartnn (hartnn):

no....since the ratio between the terms = 0.4/0.8 = 0.8/1.6 = 1/2 is common, its a geometric sequence :)